prove ln(xy)=ln(x)+ln(y)


1
Expert's answer
2021-05-24T15:20:14-0400

"\textbf{Prove:}"

Let "ln(xy)=c, ln(x)=a, ln(y)=b", then

"e^{c}=xy, \, e^{a}=x, \, e^{b}=y."

"e^{a+b}=e^{a}e^{b}=xy=e^{c}\Rightarrow a+b=c" or "ln(x)+ln(y)=ln(xy)".